hhanan:

Write the first four terms of a geometric sequence. Using the values you have written, calculate the common ratio and write the rule for the nth term. Show all calculations.

darkknight:

@supie

JammarWalker:

@darkknight

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker @darkknight \(\color{#0cbb34}{\text{End of Quote}}\) lol

darkknight:

alr well i can take this, lets make up a random geometric sequence \[An = a * r ^{n-1}\] followed by this formula

darkknight:

working backwards might be easier, lets assign r = 3 and a = 5 Now find the first 4 terms of the sequence, plug in n = 1 then 2 then 3 then 4

darkknight:

basically cheating the system, it askes u to calculate the common ratio and stuff and the eqn but we defined that ahead of time, so you can easily solve parts 2 and 3

darkknight:

just find the first 4 terms of the sequence, plug in n = 1 then 2 then 3 then 4

hhanan:

48828125

darkknight:

?

darkknight:

u should have 4 values, one for the 1st term, 1 for the second... 1 for the fourth

hhanan:

3,5,8

darkknight:

not quite... \[An = 5 * 3^{(n-1)}\] plug in 1, 2 3 and 4 for n

hhanan:

3a

darkknight:

plug in 1 for a first, what do you get for the variable (An)???

hhanan:

a

darkknight:

sorry for n, what am i saying lmao

darkknight:

plug in 1 for n,

jhonyy9:

good job @darkknight - congrats !

hhanan:

so put a1

darkknight:

\[Y = 5*3^{(n-1)}\] please plug in 1 for n and see what Y equals

hhanan:

5

darkknight:

now plug in 2

hhanan:

15

darkknight:

keep going, plug in 3 and 4

hhanan:

45,135

darkknight:

yep now ur done

darkknight:

amd thank you @jhonyy9

hhanan:

so i put all of that in for the problem

darkknight:

mhm